OpenStudy (anonymous):

help please Find the equation of the line that passes through the points (2,8) and (-5,10)

OpenStudy (anonymous):

Find the slope (first x minus first y)/(second x minus second y). Then, just put that and one point into point slope form and viola!

OpenStudy (anonymous):

ok i am doing it just a minute

OpenStudy (anonymous):

i got -3.5

OpenStudy (anonymous):

That is your slope. You still have to put it into point slope form, along with either point you wish, to get your equation.

OpenStudy (anonymous):

is point slope for y=mx+b

OpenStudy (anonymous):

is point slope form y=mx+b?

OpenStudy (anonymous):

ok is the answer y-8=-3.5(x-2)

OpenStudy (anonymous):

I believe so yes. If your teacher or whoever is picky, they may want you to put everything on one side in standard for, but that looks good.

OpenStudy (anonymous):

\[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-x1 \right)\]